OpenStudy (anonymous):

(y^2-3y+2)/(y^2+5y+6) divided by (y^2+2y-8)/(y^2+y-12)

OpenStudy (anonymous):

step1

OpenStudy (pfenn1):

\[\frac{y^2-3y+2}{y^2+5y+6} \div \frac{y^2+2y-8}{y^2+y-12}=\frac{y^2-3y+2}{y^2+5y+6}\times \frac{y^2+y-12}{y^2+2y-8}=\]

OpenStudy (anonymous):

step2

OpenStudy (pfenn1):

or you could factor each of the terms....\[\frac{y^2-3y+2}{y^2+5y+6}=\frac{(y-2)(y-1)}{(y+2)(y+3)}\]

OpenStudy (pfenn1):

\[\frac{y^2+y-12}{y^2+2y-8} = \frac{(y+4)(y-3)}{(y+4)(y-2)}\]